# Visual-Friendly Concept Protection via Selective Adversarial Perturbations

### v1

@@ Abstract.
Personalized concept generation by tuning diffusion models with a few images raises potential legal and ethical concerns regarding privacy and intellectual property rights. Researchers attempt to prevent malicious personalization using adversarial perturbations. However, previous efforts have mainly focused on the effectiveness of protection while neglecting the visibility of perturbations. They utilize global adversarial perturbations, which introduce noticeable alterations to original images and significantly degrade visual quality. In this work, we propose the Visual-Friendly Concept Protection (VCPro) framework, which prioritizes the protection of key concepts chosen by the image owner through adversarial perturbations with lower perceptibility. To ensure these perturbations are as inconspicuous as possible, we introduce a relaxed optimization objective to identify the least perceptible yet effective adversarial perturbations, solved using the Lagrangian multiplier method. Qualitative and quantitative experiments validate that VCPro achieves a better trade-off between the visibility of perturbations and protection effectiveness, effectively prioritizing the protection of target concepts in images with less perceptible perturbations.
@@
Unauthorized personalizations, adversarial examples, diffusions, privacy protections, copyright protections
copyright: acmlicensedjournal: TOMMdoi: XXXXXXX.XXXXXXXccs: Information systems Multimedia information systemsccs: Security and privacy Human and societal aspects of security and privacy
1. Introduction
@@
With the advent of popular generative models (Ho et al., 2020; Rombach et al., 2022; Song et al., 2020) such as Stable Diffusion (Rombach et al., 2022) and DALL·E 2/3 (Ramesh et al., 2022), people lacking expertise in drawing or photography can effortlessly create realistic or artistic works using simple textual descriptions. However, the success of these models has raised significant concerns about privacy, intellectual property rights, and various legal and ethical issues. For example, an adversary could easily generate sensitive specific concepts, such as personal fake images, or imitate renowned artworks for commercial purposes, using only a few reference images and concept personalization techniques such as Textual Inversion (Gal et al., 2022) or DreamBooth (Ruiz et al., 2023).
@@
Figure 1. The first row presents protected images generated by Anti-DreamBooth (Le et al., 2023), Mist (Zhang et al., 2022), MetaCloak (Liu et al., 2024), SDS(-) (Xue et al., 2024), PhotoGuard (Salman et al., 2023), and our method VCPro. The second row shows magnification sections of the background and face regions from different protected images. The third row displays the result of applying Textual Inversion to the protected data. PSNR and SSIM evaluate the visibility of adversarial perturbations; higher values indicate better results. For a fair comparison, all adversarial perturbations have a size of 8/255. Our method achieves a superior balance between final protection effectiveness and the visual quality of perturbations, prioritizing the protection of user-important concepts compared to previous works.
@@
One prevailing direction to mitigate these potential risks is to leverage adversarial attack techniques, transforming original images into adversarial examples, termed “protected images”. These protected images can misguide the personalized diffusion model and deceive its generation process, which resists malicious editing or personalization (Salman et al., 2023; Liang et al., 2023; Liang and Wu, 2023; Le et al., 2023; Shan et al., 2023; Liu et al., 2024). For instance, AdvDM (Liang et al., 2023) employs adversarial attacks in the inversion stage of Stable Diffusion to safeguard against malicious imitation of the style of a specific artist. Later, its updated version, Mist (Liang and Wu, 2023), enhances the protection efficacy of protected images by adding a textual loss, extending its application from Textual Inversion to DreamBooth. Unlike AdvDM and Mist, which focus on art style protection, Anti-DreamBooth (Le et al., 2023) undermines DreamBooth model generation quality to enhance privacy by adding adversarial perturbations to human face images before posting online. Furthermore, MetaCloak (Liu et al., 2024) leverages a meta-learning strategy and data transformations to generate more effective and robust protected images against DreamBooth.
@@
However, these methods primarily focus on preventing the personalization methods from generating high-quality corresponding images, often resulting in noticeable and unacceptable perturbations. They usually use $11/255$ (Le et al., 2023; Shan et al., 2023; Liu et al., 2024) or $17/255$ (Liang and Wu, 2023) as perturbation size in their paper, which is generally unacceptable for owners of face photos, significantly limiting the usability of these methods in real-world applications. Therefore, we pose the following question in this work: How can we find a better trade-off between visibility of perturbations and protection effectiveness?
@@
To this end, we introduce a Visual-Friendly Concept Protection (VCPro) framework to counteract unauthorized concept-driven text-to-image synthesis. This framework learns selective adversarial perturbations targeting important regions. Unlike discriminative tasks where important information is class-related and provided by target model gradients, identifying crucial information in generative tasks is challenging. VCPro utilizes user-provided masks for target concept protection, which can be supplemented by other privacy detection tools for online platforms. In that way, we propose a regional adversarial loss using spatial information to focus on selected areas. To further enhance visual quality, we apply a Lagrangian multiplier-based solution, shifting from maximizing protection effectiveness to minimizing perceptibility while ensuring effective protection. Considering human sensitivity to low-frequency changes, we measure perturbation perceptibility in the frequency domain. Experiments on models like Textual Inversion and DreamBooth validate VCPro’s effectiveness. Our approach yields subtler adversarial perturbations compared to baselines like Mist and Anti-DreamBooth, especially FID, which is reduced from $96.24$ to $27.03$.
@@ 2. Related Work
Personalization of Diffusions Models. Personalization for specific concepts (attributions, styles, or objects) has been a long-standing goal in the image generation field. In text-to-image diffusion models, previous researchers have primarily concentrated on prompt learning and test-time tuning of pre-trained models to generate images based on specific target concepts using special language tokens. Textual Inversion adjusts text embeddings of a new pseudoword to describe the concept (Gal et al., 2022). DreamBooth fine-tunes denoising networks to connect the novel concept and a less commonly used word token (Ruiz et al., 2023). Based on that, more works (Voynov et al., 2023; Zhang et al., 2023a; Kumari et al., 2023) are proposed to improve controllability and flexibility in processing image visual concepts. In this paper, we have selected Textual Inversion and DreamBooth as the techniques used by the adversary due to their popularity and representativeness.
@@
Imperceptibility Adversarial Attack. Adversarial examples (Szegedy et al., 2013; Carlini and Wagner, 2017; Duan et al., 2021; Zhang et al., 2022; Mi et al., 2023) are initially introduced by adding imperceptible noise to original data, fooling classifiers into misclassifying with high confidence. Recently, more and more researchers aim to improve the imperceptibility of adversarial examples, and they make use of a variety of tools such as perceptual color distance (Zhao et al., 2021), low-frequency spatial constraints (Luo et al., 2022), hybrid attacks in frequency and spatial domain (Jia et al., 2022), and invertible neural networks (Chen et al., 2023), etc. But they are mainly aimed at discriminative tasks such as image classification, where important information in the image can be fed back relatively accurately by the gradient of the target model, whereas we target a diffusion-based generative model whose gradient is also for the whole image.
@@
Adversarial Examples Against Unauthorized AI Generation. Unauthorized AI Generation seriously harms the safety of society and researchers are looking at several angles to reduce its risk (Zhang et al., 2024; Becattini et al., 2023; Mahto et al., 2023). Adversarial attacks are a promising way against unauthorized AI generation (Ruiz et al., 2020; Wang et al., 2022a, b; Zhu et al., 2023; Tang et al., 2024). Ruiz et al. (2020) are the first to adapt adversarial attack methods to undermine the generation of realistic deepfake images, thereby enhancing privacy security. Anti-Forgery (Wang et al., 2022a) computes adversarial noise in the Lab color space to enhance the imperceptibility and robustness of various image transformations. Unlike others, DeepFake Disrupter (Wang et al., 2022b) generates adversarial noise using an autoencoder instead of gradients and spoofs deepfake generators while remaining undetected by deepfake detectors. Zhu et al. (2023) utilize adversarial examples to protect images while simultaneously recovering original image information from tampered images. In contrast to discriminative tasks, adversarial examples play a positive role in the image generation field.
@@
With the rising popularity of diffusion models, previous works are difficult to adapt due to the obvious differences between diffusion and GAN or other generation models. Therefore, there is a growing interest in using adversarial examples to protect images from being maliciously exploited by diffusion models (Salman et al., 2023; Liang et al., 2023; Liu et al., 2024). Photoguard (Salman et al., 2023) attacks the VAE encoder or whole model to prevent images from being maliciously edited by diffusion models in the test time. At the same time, AdvDM (Liang et al., 2023) employs adversarial examples to protect artists from having their styles imitated by textual inversion by maximizing the denoising loss of a pre-trained diffusion model. As an upgraded version of AdvDM, Mist (Liang and Wu, 2023) adds a texture loss that measures the distance between the encoder representation of the original image and the perturbed image in the image encoder of latent diffusion. Similar to them, GLAZE (Shan et al., 2023) proposes to measure the art style similarity between the perturbed image and the target style using a pre-trained style-transfer model, and adversarial examples can protect the art style-related attributes in the original image directly. Different from them, Anti-DreamBooth (Le et al., 2023) protects privacy information from being learned by DreamBooth. Based on that, MetaCloak (Liu et al., 2024) uses a meta-learning strategy and data transformations to generate better effective and robust adversarial examples. Meanwhile, Xue et al. (2024) introduce Score Distillation Sampling (SDS) loss into the perturbation optimization process to reduce computation cost and find maximize or minimize SDS loss both achieve competitive protection effects.
@@
3. Preliminaries
@@
(1) $\mathcal{L_{\theta}}:=\mathbb{E}_{z\sim\mathcal{E}(x),y,g\sim\mathcal{N}(0,1),t}\left[\left\|{g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right)}\right\|_{2}^{2}\right],$
@@
Based on LDM, Textual Inversion, and DreamBooth are two main techniques for personalization. Textual Inversion learns new special tokens containing specific concepts to store new knowledge from images. In Textual Inversion, a pseudo token $sks$ and its corresponding embedding vector $v\in\mathcal{P}$ are introduced to describe the concept, where $\mathcal{P}$ is the word embedding space of the text encoder. They optimize $v$ directly by minimizing the objective specified in Eq. (1)s using a small set of images representing the concept as training data.
(2) $\arg\min_{v}\mathbb{E}_{z,y,g,t}\left[\left\|g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y,v)\right)\right\|_{2}^{2}\right].$
@@
(3) $\begin{gathered}\theta:=\arg\min_{\theta}\mathbb{E}_{z,y,g,t}\Bigg{[}\left\|g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y,v)\right)\right\|_{2}^{2}+\left\|g-\varepsilon_{\theta}\left(z^{p}_{t},t^{p},\tau_{\theta}(y^{p})\right)\right\|_{2}^{2}\Bigg{]}.\end{gathered}$
Protected image for Diffusions. Protected images (adversarial examples) aim to add an imperceptible adversarial perturbation $\delta$ to the original image $x$ to make the full process of condition generations ineffective. The protected image $\mathcal{L_{\theta}}$ can be formalized as:
(4) $\begin{gathered}x^{\prime}:=arg\max_{x^{\prime}}\mathcal{L_{\theta}}(x^{\prime},y),\\ \text{ s.t. }\quad\left\|x-x^{\prime}\right\|\leq\epsilon,\end{gathered}$
@@
4. Visual-Friendly Concept Protection
4.1. Overview
Fig. 2 shows the pipeline of the proposed framework for visual-friendly concept protection. Accurately describing spatial positions through language can be challenging for users, but it can be precisely achieved using masks. By leveraging SAM (Kirillov et al., 2023) or other segmentation tools, users can generate a mask $m$ for important concepts within a given image $x$. The user-provided masks and original images are then collectively fed into the protected image generation module as described in Eq. (8). In this module, we propose a regional adversarial learning loss to reduce the visibility of protected images through precise protection and a Lagrangian multiplier-based solution to minimize perturbations while maintaining successful protection.
In this section, we start by formulating the regional adversarial learning framework for diffusion models in Sec. 4.2 and then move on to the solution of the proposed optimization objectives in Sec. 4.3. The total learning process is shown in Alg. 1.
Figure 2. The pipeline of VCPro. Initially, users can utilize SAM or alternative segmentation tools to create user-protective masks within images. Following this, the masks and the original images are simultaneously input into the protected image generation module, as described by Eq. (8). Within this module, we introduce a regional adversarial learning loss aimed at diminishing the visibility of protected images by precise protection and transform objectives to minimize perturbations while maintaining effective protection.
4.2. Formulation
@@
(5) $\begin{gathered}x^{\prime}:=\arg\max_{x^{\prime}}\mathcal{L}^{\prime}_{\theta}\left(x^{\prime},y,m\right),\\ \text{ s.t. }\quad\left\|x-x^{\prime}\right\|\leq\epsilon,\\ \end{gathered}$
and the regional adversarial loss $\mathcal{L}^{\prime}_{\theta}$ is :
(6) $\mathcal{L}^{\prime}_{\theta}:=$ $\mathbb{E}_{z\sim\mathcal{E}(x),y,g\sim\mathcal{N}(0,1),t}\left[l_{mask}(z_{t},y,g,t,m)\right],$
$\text{where }\ l_{mask}(z_{t},y,g,t,m):=$ $\left\|g\odot m-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y,v)\right)\odot m\right\|_{2}^{2}-\left\|g\odot(1-m)-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y,v)\right)\odot(1-m)\right\|_{2}^{2}.$
In that way, protected image $x^{\prime}$ pushes the output of the diffusion models far away from the ground truth Gaussian noise $g$ within a specified mask region, while predictions in other areas pull close. Without the pull component of the regional adversarial loss, the part outside the mask can not provide gradient feedback, making it difficult to optimize. It affects the optimization against perturbations, so we add the pull module, and removing the pull module will affect the protection effect as shown in Table 3.
4.3. Lagrangian Multiplier-based Loose Solution
Based on Eq. (6), protected images will destroy the target area as much as possible. However, for the generation task, the output can still be recognized as synthetic as long as there are some clear signs of protection. Typical image classification attacking methods (Carlini and Wagner, 2017; Luo et al., 2022; Chen et al., 2023) also present similar viewpoints that minimize the impact on normal visual perception while maintaining the effectiveness of the attack. Hence, we propose a loosed optimization objective by attempting to find the minimal adversarial perturbation $\delta$ that can attack diffusion models successfully. For convenience, we set a heuristic method: if $\mathcal{L}^{\prime}_{\theta}>\alpha$, the attack is successful. The problem can be translated as:
(7) minimize $D(x,x+\delta),$
such that $-\mathcal{L}^{\prime}_{\theta}+\alpha\leq 0,$
$\delta\in[-\epsilon,\epsilon]^{n},$
$x+\delta\in[0,255]^{n},$
where $D(\cdot)$ is a distance metric.
(8) minimize $\mathcal{L}_{final}=D(x,x+\delta)+c\cdot(-\mathcal{L}^{\prime}_{\theta}+\alpha),$
such that $\delta\in[-\epsilon,\epsilon]^{n},$
$x+\delta\in[0,255]^{n},$
where a constant $c>0$ is appropriately selected. The equivalence of the Eq. (7) and Eq. (8) can be understood by the existence of a positive constant $c$ ensuring the best solution for the second formulation aligns with that of the first.
Considering that the human visual system is more sensitive to low-frequency regions (Luo et al., 2022), we use $D(\cdot)$ to limit perturbations into the high-frequency regions. Specifically, we use discrete wavelet transform (DWT) to transform images from the spatial domain to the frequency domain. DWT will decompose the image $x$ into one low-frequency and three high-frequency components, i.e., $x_{ll},x_{lh},x_{hl},x_{hh}$, and inverse DWT (IDWT) uses all four components to reconstruct the image.
$x_{ll}=LxL^{T},\quad x_{lh}=HxL^{T},$
$x_{hl}=LxH^{T},\quad x_{hh}=HxH^{T},$
@@ 4.3. Lagrangian Multiplier-based Loose Solution
5. Experiments

### v2

@@ Abstract
Personalized concept generation by tuning diffusion models with a few images raises potential legal and ethical concerns regarding privacy and intellectual property rights. Researchers attempt to prevent malicious personalization using adversarial perturbations. However, previous efforts have mainly focused on the effectiveness of protection while neglecting the visibility of perturbations. They utilize global adversarial perturbations, which introduce noticeable alterations to original images and significantly degrade visual quality. In this work, we propose the Visual-Friendly Concept Protection (VCPro) framework, which prioritizes the protection of key concepts chosen by the image owner through adversarial perturbations with lower perceptibility. To ensure these perturbations are as inconspicuous as possible, we introduce a relaxed optimization objective to identify the least perceptible yet effective adversarial perturbations, solved using the Lagrangian multiplier method. Qualitative and quantitative experiments validate that VCPro achieves a better trade-off between the visibility of perturbations and protection effectiveness, effectively prioritizing the protection of target concepts in images with less perceptible perturbations. Our code is available at https://github.com/KululuMi/VCPro.
Index Terms:
@@
Unauthorized personalizations, Adversarial Attacks, Diffusions, Privacy protections, Copyright protections.
I Introduction
@@
With the advent of popular image generative models [1, 2, 3] such as Stable Diffusion [2] and GPT-4o [4], people lacking expertise in drawing or photography can effortlessly create realistic or artistic works using simple textual descriptions. However, the success of these models has raised significant concerns about privacy, intellectual property rights, and various legal and ethical issues [5]. For example, an adversary could easily generate sensitive specific concepts, such as personal fake images, or imitate renowned artworks for commercial purposes, using only a few reference images and concept personalization techniques such as Textual Inversion [6] or DreamBooth [7].
@@
Figure 1: The first row shows protected images from Anti-DreamBooth [8], Mist [9], MetaCloak [10], SDS(-) [11], PhotoGuard [12], and our VCPro. The second row highlights magnified background and face regions, while the third row presents Textual Inversion results. With 8/255 perturbation size, our method achieves a better balance between protection effectiveness and visual quality for user-important concepts, as shown by higher PSNR/SSIM values.
@@
One prevailing direction to mitigate these potential risks is to leverage adversarial attack techniques, transforming original images into adversarial examples, termed “protected images”. These protected images can misguide the personalized diffusion model and deceive its generation process, which resists malicious editing or personalization [12, 13, 14, 8, 15, 10]. For instance, AdvDM [13] employs adversarial attacks in the inversion stage of Stable Diffusion to safeguard against malicious imitation of the style of a specific artist. Later, its updated version, Mist [14], enhances the protection efficacy of protected images by adding a textual loss, extending its application from Textual Inversion to DreamBooth. Unlike AdvDM and Mist, which focus on art style protection, Anti-DreamBooth [8] undermines DreamBooth model generation quality to enhance privacy by adding adversarial perturbations to human face images before posting online. Furthermore, MetaCloak [10] leverages a meta-learning strategy and data transformations to generate more effective and robust protected images against DreamBooth.
@@
However, these methods primarily focus on preventing the personalization methods from generating high-quality corresponding images, often resulting in noticeable and unacceptable perturbations. They usually use $11/255$ [8, 15, 10] or $17/255$ [14] as perturbation size in their paper, which is generally unacceptable for owners of face photos, significantly limiting the usability of these methods in real-world applications. Therefore, we pose the following question in this work: How can we find a better trade-off between visibility of perturbations and protection effectiveness?
@@
To this end, we introduce a Visual-Friendly Concept Protection (VCPro) framework to counteract unauthorized concept-driven text-to-image synthesis. This framework learns selective adversarial perturbations targeting important regions. Unlike discriminative tasks where important information is class-related and provided by target model gradients, identifying crucial information in generative tasks is challenging. VCPro utilizes user-provided masks for target concept protection, which can be supplemented by other privacy detection tools for online platforms. In that way, we propose a regional adversarial loss using spatial information to focus on selected areas. To further enhance visual quality, we apply a Lagrangian multiplier-based solution, shifting from maximizing protection effectiveness to minimizing perceptibility while ensuring effective protection. Considering human sensitivity to low-frequency changes, we measure perturbation perceptibility in the frequency domain. Experiments on models like Textual Inversion and DreamBooth validate VCPro’s effectiveness. Our approach yields subtler adversarial perturbations compared to baselines like Mist and Anti-DreamBooth, especially FID, which is reduced from $96.24$ to $27.04$.
@@ II Related Work
Personalization of Diffusions Models. Personalization for specific concepts (attributions, styles, or objects) has been a long-standing goal in the image generation field. In text-to-image diffusion models, previous researchers have primarily concentrated on prompt learning and test-time tuning of pre-trained models to generate images based on specific target concepts using special language tokens. Textual Inversion adjusts text embeddings of a new pseudoword to describe the concept [6]. DreamBooth fine-tunes denoising networks to connect the novel concept and a less commonly used word token [7]. Based on that, more works [16, 17, 18] are proposed to improve controllability and flexibility in processing image visual concepts. In this paper, we have selected Textual Inversion and DreamBooth as the techniques used by the adversary due to their popularity and representativeness.
@@
Imperceptibility Adversarial Attack. Adversarial examples [19, 20, 21, 9, 22] are initially introduced by adding imperceptible noise to original data, fooling classifiers into misclassifying with high confidence. Recently, more and more researchers aim to improve the imperceptibility of adversarial examples, and they make use of a variety of tools such as perceptual color distance [23], low-frequency spatial constraints [24], hybrid attacks in frequency and spatial domain [25], and invertible neural networks [26], etc. But they are mainly aimed at discriminative tasks such as image classification, where important information in the image can be fed back relatively accurately by the gradient of the target model, whereas we target a diffusion-based generative model whose gradient is also for the whole image.
@@
Adversarial Examples Against Unauthorized AI Generation. Unauthorized AI Generation seriously harms the safety of society, prompting researchers to explore multiple approaches for mitigating these risks. While passive defense methods focus on distinguishing between authentic and synthetic images [27, 28], adversarial attacks are a promising way against unauthorized AI generation [29, 30, 31, 32]. Ruiz et al. [29] are the first to adapt adversarial attack methods to undermine the generation of realistic deepfake images, thereby enhancing privacy security. Anti-Forgery [30] computes adversarial noise in the Lab color space to enhance the imperceptibility and robustness of various image transformations. Unlike others, DeepFake Disrupter [31] generates adversarial noise using an autoencoder instead of gradients and spoofs deepfake generators while remaining undetected by deepfake detectors. Zhu et al. [32] utilize adversarial examples to protect images while simultaneously recovering original image information from tampered images. In contrast to discriminative tasks, adversarial examples play a positive role in the image generation field.
@@
With the rising popularity of diffusion models, previous works are difficult to adapt due to the obvious differences between diffusion and GAN or other generation models. Therefore, there is a growing interest in using adversarial examples to protect images from being maliciously exploited by diffusion models [12, 13, 10]. Photoguard [12] attacks the VAE encoder or whole model to prevent images from being maliciously edited by diffusion models at the test time. At the same time, AdvDM [13] employs adversarial examples to protect artists from having their styles imitated by textual inversion by maximizing the denoising loss of a pre-trained diffusion model. As an upgraded version of AdvDM, Mist [14] adds a texture loss that measures the distance between the encoder representation of the original image and the perturbed image in the image encoder of latent diffusion. Similar to them, GLAZE [15] proposes to measure the art style similarity between the perturbed image and the target style using a pre-trained style-transfer model, and adversarial examples can protect the art style-related attributes in the original image directly. Different from them, Anti-DreamBooth [8] protects privacy information from being learned by DreamBooth. Based on that, MetaCloak [10] uses a meta-learning strategy and data transformations to generate better effective and robust adversarial examples. Meanwhile, Xue et al. [11] introduce Score Distillation Sampling (SDS) loss into the perturbation optimization process to reduce computation cost and find maximize or minimize SDS loss both achieve competitive protection effects.
@@
Figure 2: The pipeline of VCPro. Initially, users can utilize SAM or alternative segmentation tools to create user-protective masks within images. Following this, the masks and the original images are simultaneously input into the protected image generation module, as described by Eq. (8). Within this module, we introduce a regional adversarial learning loss aimed at diminishing the visibility of protected images by precise protection and transform objectives to minimize perturbations while maintaining effective protection.
III Preliminaries
@@
$\mathcal{L_{\theta}}:=\mathbb{E}_{z\sim\mathcal{E}(x),y,g\sim\mathcal{N}(0,1),t}\left[\left\|{g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right)}\right\|_{2}^{2}\right],$ (1)
@@
Based on LDM, Textual Inversion, and DreamBooth are two main techniques for personalization. Textual Inversion learns new special tokens containing specific concepts to store new knowledge from images. In Textual Inversion, $y$ usually is a generic prompt, such as “a photo of sks [class noun]”, with $sks$ is a pseudo token, and “[class noun]” is the object type. They optimize the embedding of $sks$ directly by minimizing the objective specified in Eq. (1), which uses a small set of images representing the concept as training data.
$\arg\min_{v}\mathbb{E}_{z,y,g,t}\left[\left\|g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right)\right\|_{2}^{2}\right].$ (2)
@@
$\begin{gathered}\theta:=\arg\min_{\theta}\mathbb{E}_{z,y,g,t}\Bigg{[}\left\|g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right)\right\|_{2}^{2}\\ +\left\|g-\varepsilon_{\theta}\left(z^{p}_{t},t^{p},\tau_{\theta}(y^{p})\right)\right\|_{2}^{2}\Bigg{]}.\end{gathered}$ (3)
Protected Image for Diffusions. Protected images (adversarial examples) aim to add an imperceptible adversarial perturbation $\delta$ to the original image $x$ to make the full process of condition generation ineffective. The protected image $\mathcal{L_{\theta}}$ can be formalized as:
$\begin{gathered}x^{\prime}:=arg\max_{x^{\prime}}\mathcal{L_{\theta}}(x^{\prime},y),\\ \text{ s.t. }\quad\left\|x-x^{\prime}\right\|\leq\epsilon,\end{gathered}$ (4)
@@
IV Visual-Friendly Concept Protection
IV-A Overview
Fig. 2 shows the pipeline of the proposed framework for visual-friendly concept protection. Accurately describing spatial positions through language can be challenging for users, but it can be precisely achieved using masks. By leveraging SAM [33] or other segmentation tools, users can generate a mask $m$ for important concepts within a given image $x$. The user-provided masks and original images are then collectively fed into the protected image generation module as described in Eq. (8). In this module, we propose a regional adversarial learning loss to reduce the visibility of protected images through precise protection and a Lagrangian multiplier-based solution to minimize perturbations while maintaining successful protection.
In this section, we start by formulating the regional adversarial learning framework for diffusion models in Sec. IV-B and then move on to the solution of the proposed optimization objectives in Sec. IV-C. The total learning process is shown in Alg. 1.
IV-B Formulation
@@
$\begin{gathered}x^{\prime}:=\arg\max_{x^{\prime}}\mathcal{L}^{\prime}_{\theta}\left(x^{\prime},y,m\right),\\ \text{ s.t. }\quad\left\|x-x^{\prime}\right\|\leq\epsilon,\\ \end{gathered}$ (5)
and the regional adversarial loss $\mathcal{L}^{\prime}_{\theta}$ combines a “push” term for protected regions and a “pull” term for non-protected regions.
$\mathcal{L}^{\prime}_{\theta}:=$ $\mathbb{E}_{z\sim\mathcal{E}(x),y,g\sim\mathcal{N}(0,1),t}\left[l_{mask}(z_{t},y,g,t,m)\right],$ (6)
$l_{mask}(z_{t},y,g,t,m):=$ $\underbrace{\left\|(g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right))\odot m\right\|_{2}^{2}}_{\text{push term for protected regions}}$
$-\underbrace{\left\|(g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right))\odot(1-m)\right\|_{2}^{2}}_{\text{pull term for non-protected regions}}.$
This loss operates through a balanced mechanism of opposing forces. The push term maximizes the distance between the predicted noise $\varepsilon_{\theta}$ and ground truth noise $g$ in masked regions ($m=1$), effectively disrupting the denoising process for protected concepts. Simultaneously, the pull term minimizes this distance in unmasked regions ($m=0$), preserving visual quality in non-protected areas.
During optimization, these components generate distinct gradient signals: push gradients divert predictions away from ground truth in protected regions, while pull gradients maintain accuracy elsewhere. This dual approach is crucial for effective protection—without the pull component, regions outside the mask cannot provide sufficient gradient feedback, significantly impairing the optimization process. Our ablation studies in Fig. 4 and Table II demonstrate that removing the pull component substantially degrades protection effectiveness.
@@
IV-C Lagrangian Multiplier-based Loose Solution
Based on Eq. (6), protected images will destroy the target area as much as possible. However, for the generation task, the output can still be recognized as synthetic as long as there are some clear signs of protection. Typical image classification attacking methods [20, 24, 26] also present similar viewpoints that minimize the impact on normal visual perception while maintaining the effectiveness of the attack. Hence, we propose a loosed optimization objective by attempting to find the minimal adversarial perturbation $\delta$ that can attack diffusion models successfully. For convenience, we set a heuristic method: if $\mathcal{L}^{\prime}_{\theta}>\alpha$, the attack is successful. The problem can be translated as:
minimize $D(x,x+\delta),$ (7)
such that $-\mathcal{L}^{\prime}_{\theta}+\alpha\leq 0,$
$\delta\in[-\epsilon,\epsilon]^{n},$
$x+\delta\in[0,255]^{n},$
where $D(\cdot)$ is a distance metric.
minimize $\mathcal{L}_{final}=c\cdot D(x,x+\delta)-\mathcal{L}^{\prime}_{\theta}+\alpha,$ (8)
such that $\delta\in[-\epsilon,\epsilon]^{n},$
$x+\delta\in[0,255]^{n},$
where a constant $c>0$ is appropriately selected. The equivalence of Eq. (7) and Eq. (8) can be understood by the existence of a positive constant $c$ ensuring the best solution for the second formulation aligns with that of the first.
Considering that the human visual system is more sensitive to low-frequency regions [24], we use $D(\cdot)$ to limit perturbations into the high-frequency regions. Specifically, we use discrete wavelet transform (DWT) to transform images from the spatial domain to the frequency domain. DWT will decompose the image $x$ into one low-frequency and three high-frequency components, i.e., $x_{ll},x_{lh},x_{hl},x_{hh}$, and inverse DWT (IDWT) uses all four components to reconstruct the image.
$x_{ll}=LxL^{T},\quad x_{lh}=HxL^{T},$
$x_{hl}=LxH^{T},\quad x_{hh}=HxH^{T},$
V Experiments